Topics In Contemporary Math
Exponential Growth
4) A culture of paramecia contains 75 organisms. Assuming the presence of sufficient food
and physical room, the culture increases by 30% per hour. How many paramecia would
there be after 12 hours?

Found 2 solutions by math_tutor2020, Alan3354:
Answer by math_tutor2020(3817)


x = number of hours
y = population count

The exponential function that models this scenario is
y = 75*(1.30)^x
It's in the form
y = a*b^x

a = 75 = initial population
b = 1.30 = growth multiplier
The 1.30 represents 30% growth

Plug in x = 12 to find the following
y = 75*(1.30)^x
y = 75*(1.30)^12
y = 1747.35638418608
y = 1747

Answer: approximately 1747


Answer by Alan3354(69443)

A culture of paramecia contains 75 organisms. Assuming the presence of sufficient food and physical room, the culture increases by 30% per hour. How many paramecia would there be after 12 hours?

---------------------
75*1.3^12